Find an expression for the function whose graph consists of the line segment from the point (-2,2) to the point (-1,0) together with the top half of the circle with centere at the origin and radius 1.
1
Expert's answer
2020-04-16T18:54:00-0400

Equation of line joining the two points "(-2,2)" and "(-1,0)" is,



"y-2=\frac{0-2}{-1+2}(x+2)"



"y-2=-2(x+2)"



"y=-2x-2"


Here, the line segment lies in the interval "[-2,-1]".


Now, consider the circle of radius "1" centered at origin as "x^2+y^2=1"


So, the top half of the circle is given by,



"y=\sqrt{1-x^2}"

Here, the circular arc lies in the interval "[-1,1]".


Therefore, the function is defined piece-wise and the expression for the function is given as piece-wise function,

"f(x) = \begin{cases}
 -2x-2 &\text{if } -2\le x\le -1\\
 \sqrt{1-x^2} &\text{if } -1\le x\le 1
\end{cases}"

The sketch of the curve is as shown in the figure below: